# Confidential Prompting: Protecting User Prompts from Cloud LLM Providers

### v2

@@ 6 Implementation
We implement OSMD (SMD+PO) based on PyTorch and target the widely adopted Llama [19] series of open-source LLMs. We next describe a few key implementations in our prototype.
LLM Serving Server.  We develop an LLM serving server based on the Transformers library [45], which assumes that all user requests arrive simultaneously (no continuous batching), for simplicity of implementation. To support SMD, we adapted the Llama model by monkey-patching its attention module. Specifically, we modified the attention score calculation to asynchronously send the query tensor ($Q$ in Theorem 4.1) to each user via the GLOO tensor communication backend [46]. This allows us to receive private attention scores from users while computing public attention scores in parallel. The server waits for all private scores to arrive before evaluating the final attention score.
Prompt Vault.  Within each user’s CVM, operations like prefill, PO, and calculating private attention scores are managed by a client called the prompt vault. The prompt vault synchronously listens to the LLM serving server to receive the $Q$ tensor over the GLOO backend and returns results based on Theorem 4.1.
Serving virtual prompts with sub-linear memory.  We implement the compact representation mechanism for virtual prompts (§5.2) using asymmetric attention masks. For example, suppose we have two prompts—real and fake—that share a common segment $S_{\text{shared}}$ of length $a$. The real token sequence’s KV cache is $S_{\text{real}}$ , and the fake one is $S_{\text{fake}}$ , both of length $b$. The naive way to store attention states is to batch them:
@@ 7 Evaluation
Figure 7: $(\epsilon,\lambda)$ pairs sampled from the clinical dataset with $\lambda_{\text{max}}=512$, for locational PII, e.g. address and name of places. Unless otherwise specified, the sampling temperature is set to $\tau$ = 1.
@@
Figure 12: Overhead breakdown of SMD in the 13B model. As the number of users increases, the communication overhead of sending $Q$ and receiving private attention increases linearly.
@@
  • Does our design preserve user prompt confidentiality in practice?
  • Does our design meet the compute efficiency requirement, considering that output invariance and model confidentiality for SMD and PO are trivial?
To answer the first question, we empirically examine the maximum available $\lambda$ for the desired error bounds given by $\epsilon$, which are referred to as pairs of security factors $(\epsilon,\lambda)$ in the following, and examine other factors that influence the sampling results. Subsequently, we evaluate the performance of the SMD + PO method compared to two naive methods based on CVM as presented in Figure 2.
Evaluation setup.  Our evaluation utilizes the Llama 2 [19] models with 7B and 13B parameters, Llama 3 [48] model with 8B parameters, and Code Llama [49] model with 34B parameters. All experiments were conducted on a cloud node equipped with an NVIDIA H100 NVL GPU with 94 GB of memory, 40 AMD EPYC Genoa processor cores, and 320 GB of system memory. Architectural security features such as AMD SEV-SNP [25] and NVIDIA GPU CC [27] are enabled for evaluation that require confidential computing. It is worth noting that our design does not require confidential computing protection for the LLM serving server. However, since computing nodes with multiple GPUs that support enabling GPU CC feature were not available at the time of writing, and because NVIDIA GPU CC feature can only be enabled or disabled entirely on a single GPU, the confidential computing overhead also applies to the LLM server in the following experiments. Consequently, the performance overhead presented in this section is higher than the ideal case described in our design.
@@ 7.1 Security of Prompt Obfuscation
Effect of Different Models and Parameter Size.  The impact of the model parameter size on $\lambda$ is relatively small. For example, comparing Llama 2 models with 7B and 13B parameters on the clinical dataset results in only a 4% difference in $\lambda$. In contrast, the choice of language model has a much more substantial effect on $\lambda$. As shown in Figure 7, there is a 2.4$\times$ difference in $\lambda$ between Llama 3 and Llama 2 (13B). We observed that Llama 3, as a more advanced model than Llama 2, is more “confident” in selecting the next token, which narrows the possible candidate pool. In other words, the distribution modeled by Llama 2 is flatter, resulting in a higher $\lambda$ under the same conditions.
@@ 7.2 Performance Analysis
We implement two baselines for performance evaluation to demonstrate the scalability and compute efficiency of our approach: (1) No protection, a naive approach where every user is served with a single LLM instance within a single CVM (2(b)), which is intended to demonstrate the upper bound for the performance with confidential computing; and (2) Full isolation, a naive approach where each user is served with a dedicated LLM instance within a per-user CVM (2(c)). Our approach, which uses SMD (2(d)), is denoted as SMD. We compare the performance of our approach with that of Full isolation because the latter defends against a similar threat model as ours.
@@ 7.2.1 Scalability
Number of Users.  The primary challenge in using CVMs for LLM services is achieving scalability with a growing number of users (§2.2). Full isolation handle separate instances of the LLM for each user, which limits the maximum number of concurrent user CVMs due to GPU memory limitations. As shown in Figure 9, the latency of Full isolation increases significantly as the number of users increases. In contrast, the SMD approach scales effectively with the number of users, because it can host a large number of user CVMs concurrently.
Model Parameter Size.  The Full isolation baseline also faces scalability challenges as model parameter sizes increase. In contrast, SMD is less affected by parameter size scaling. As presented in Figure 9, Full isolation’s end-to-end latency scales at a higher rate than that of SMD under the same conditions, when the model size increases from 7B to 34B.
Number of Input/Output Tokens.  The number of input tokens primarily impacts the initial time-to-first-token (TTFT) latency. Both Full isolation and SMD experience a 15$\times$ higher TTFT latency compared to the No protection baseline due to the absence of batch parallelism during the prefill phase. In contrast, the number of input tokens has negligible impact on the end-to-end latency in SMD, which reflects the time of decoding. As shown in Figure 10, the three bars in each group represent end-to-end latency for different input token counts such as 64, 128 and 256, with negligible differences among them, across different combinations of output token count and $\lambda$. Meanwhile, the increase of output token count also has negligible impact on latency per token, as decoding each token has static overhead.
@@ 7.2.2 Compute efficiency
We further analyze the overhead of SMD and PO, along with their breakdown, to demonstrate that our design is computationally efficient.
Overhead of Secure Multi-party Decoding.  Secure Multi-party Decoding introduces overhead due to (1) confidential computation (CC), (2) the absence of batch processing in CVMs, and (3) communication between CVMs and the LLM. For (1), Figure 12 illustrates the impact of confidential computing (CC) on the No protection baseline and SMD. It shows that CC results in approximately $15\%$ overhead compared to tranditional computing without CC. For (2) and (3), we present SMD’s latency breakdown in Figure 12. The six latency sources align with the processing steps in Figure 3. The QKV projection, public attention computation and MLP are common in traditional decoding. However, the unique overhead introduced by SMD stems from the other three sources. Notably, the “Receiving private attention” latency includes both the time the cloud LLM spends waiting for private attention computation to finish and the communication cost. This encapsulates the overhead from (2) and (3), as either the absence of batch processing or high communication costs can cause the LLM to remain idle. As shown in Figure 12, the latency of (2) and (3) (“Sending $Q_{new}$” and “Receiving private attention”) dominates the overhead, accounting for up to $40\%$ of the total latency. In contrast, the overhead from merging attentions is relatively minor.
Overhead of Prompt Obfuscation.  Prompt Obfuscation introduces overhead due to (1) sampling fake n-grams for virtual prompts, which increases TTFT, and (2) decoding for $\lambda$ virtual prompts. For (1), Figure 14 illustrates the sampling latency for virtual prompts. The average sampling time for an eight-token replacement using $\lambda=512$ and $\epsilon=1/32$ is around 1 second. Sampling multiple subsequences can be batched, for example, 16 subsequences take about 1.5 seconds. Lower $\epsilon$ results in faster sampling due to a reduced $\lambda$. For (2), the impact of decoding virtual prompts on latency is shown in Figure 14. When $\lambda\leq 8$, the overhead is less than $3\%$. However, the overhead increases as $\lambda$ grows: at $\lambda=64$ and $\lambda=128$, the overhead is approximately $7\%$ and $30\%$, respectively. This overhead also scales with the number of users. As shown in Figure 10, for 8 users, the overhead of $\lambda=8$ and $\lambda=16$ is around $8\%$ and $25\%$, respectively. Fundamentally, the overhead grows with the total number of prompts increases, as the LLM processes user prompts and virtual prompts equally. On the other hand, virtual prompts impose minimal overhead on the CVM memory footprint, thanks to our memory-efficient approach for storing virtual prompts (§5.2). In conclusion, the use of Prompt Obfuscation and its associated overhead reflects a key trade-off in our design, as outlined in §3.1, that is balancing prompt confidentiality (§5.3) with compute efficiency.
@@ 8 Concluding Remarks
Cloud-based LLM services are becoming pervasive in our daily lives, but raise privacy concerns since users must submit their prompts to the cloud. OSMD combines two techniques to protect user prompts from LLM providers: secure multiparty decoding (SMD) and prompt obfuscation (PO). SMD fully leverages the confidential computing capabilities of modern hardware for efficient and scalable confidential prompting. PO, on the other hand, leverages the statistical properties of the LLM itself to achieve confidentiality without encryption. Our proposed techniques have the potential to enable privacy-preserving LLM applications such as chatbots and AI assistants that interact with a large number of end users, without high operational costs. Next, we discuss some limitations in OSMD as well as new opportunities.

### v3

@@ 6 Implementation
We implement OSPD (SPD+PO) based on PyTorch and target the widely adopted Llama [55] series of open-source LLMs. We next describe a few key implementations in our prototype.
LLM Serving Server.  We develop an LLM serving server based on the Transformers library [58], which assumes all user requests arrive simultaneously (no continuous batching) for simplicity. To support SPD, we adapted the Llama model by monkey-patching its attention module. We modified the attention score calculation to asynchronously send the query tensor ($Q$ in Theorem 4.1) to each user via the GLOO tensor communication backend [16]. This allows us to receive private attention scores from users while computing public attention scores in parallel. The server waits for all private scores to arrive before evaluating the final attention score.
Prompt Vault.  Within each user’s CVM, a client called the prompt vault manages operations like prefill, PO, and calculating private attention scores. The prompt vault synchronously listens to the LLM serving server to receive the $Q$ tensor over the GLOO backend and returns results based on Theorem 4.1.
Serving virtual prompts with sub-linear memory.  We implement the compact representation mechanism for virtual prompts (§5.1) using asymmetric attention masks. For example, suppose we have two prompts—real and fake—sharing a common segment $S_{\text{shared}}$ of length $a$. The real sequence’s KV cache is $S_{\text{real}}$ , and the fake one is $S_{\text{fake}}$ , both of length $b$. The naive way to store attention states is to batch them:
@@ 7 Evaluation
Figure 6: $(\epsilon,\lambda)$ pairs sampled from the clinical dataset with $\lambda_{\text{max}}=512$, for locational PII, e.g. name of places. Unless otherwise specified, the sampling temperature is set to $\tau$ = 1.
@@
Figure 11: Overhead breakdown of SPD in the 13B model. As the number of users increases, the communication overhead of sending $Q$ and receiving private attention increases linearly.
@@
  • Does OSPD preserve prompt confidentiality in practice?
  • Does OSPD maintain compute efficiency, considering model confidentiality and output invariance are intuitive?
To answer the first question, we empirically examine the maximum available $\lambda$ for the desired error bounds given by $\epsilon$, which are referred to as pairs of security factors $(\epsilon,\lambda)$ in the following, and examine other factors that influence the sampling results. Subsequently, we evaluate the performance of OSPD compared to two naive CVM methods (Figure 2).
Evaluation setup.  Our evaluation utilizes the Llama 2 [55] models with 7B and 13B parameters, Llama 3 [35] model with 8B parameters, and Code Llama [34] model with 34B parameters. All experiments were conducted on a cloud node equipped with an NVIDIA H100 NVL GPU with 94 GB of memory, 40 AMD EPYC Genoa processor cores, and 320 GB of system memory. Architectural security features such as AMD SEV-SNP [3] and NVIDIA GPU CC [40] are enabled for evaluation that require confidential computing. We note that our design does not require confidential computing protection for the LLM serving server. However, since computing nodes with multiple GPUs that support enabling GPU CC feature were not available at the time of writing, and because NVIDIA GPU CC feature can only be enabled or disabled entirely on a single GPU, the confidential computing overhead also applies to the LLM server in the following experiments. Consequently, the performance overhead presented in this section is higher than the ideal case described in our design.
@@ 7.1 Security of Prompt Obfuscation
Effect of Different Models and Parameter Size.  The impact of the model parameter size on $\lambda$ is relatively small. For example, comparing Llama 2 models with 7B and 13B parameters on the clinical dataset results in only a 4% difference in $\lambda$. In contrast, the choice of language model has a more substantial effect on $\lambda$. In Figure 6, there is a 2.4$\times$ difference in $\lambda$ between Llama 3 and Llama 2 (13B). We observed that Llama 3, as a more advanced model than Llama 2, is more “confident” in selecting the next token, which narrows the possible candidate pool. In other words, the distribution modeled by Llama 2 is flatter, resulting in a higher $\lambda$ under the same conditions.
@@ 7.2 Performance Analysis
We implement two baselines for performance evaluation to demonstrate the scalability and compute efficiency of our approach: (1) No protection, a naive approach where every user is served with a LLM instance within a CVM (2(b)), which is intended to demonstrate the upper bound for the performance with confidential computing; and (2) Full isolation, a naive approach where each user is served with a dedicated LLM instance in a per-user CVM (2(c)). Our approach is denoted as SPD (2(d)). We compare the performance of our approach with that of Full isolation because the latter defends against a similar threat model as ours.
@@ 7.2.1 Scalability
Number of Users.  The primary challenge in using CVMs for LLM services is achieving scalability with a growing number of users (§2.2). Full isolation handle separate instances of the LLM for each user, which limits the maximum number of concurrent user CVMs due to GPU memory limitations. As shown in Figure 8, the latency of Full isolation increases significantly as the number of users increases. In contrast, the SPD approach scales effectively with the number of users, because it can host a large number of user CVMs concurrently.
Model Parameter Size.  The Full isolation baseline also faces scalability challenges as model parameter sizes increase. In contrast, SPD is less affected by parameter size scaling. As presented in Figure 8, Full isolation’s end-to-end latency scales at a higher rate than that of SPD under the same conditions, when the model size increases from 7B to 34B.
Number of Input/Output Tokens.  The number of input tokens primarily impacts the initial time-to-first-token (TTFT) latency. Both Full isolation and SPD experience a 15$\times$ higher TTFT latency compared to the No protection baseline due to the absence of batch parallelism during the prefill phase. In contrast, the number of input tokens has negligible impact on the end-to-end latency in SPD, which reflects the time of decoding. As shown in Figure 9, the three bars in each group represent end-to-end latency for different input token counts such as 64, 128 and 256, with negligible differences among them, across different combinations of output token count and $\lambda$. Meanwhile, the increase of output token count also has negligible impact on latency per token, as decoding each token has static overhead.
@@ 7.2.2 Compute efficiency
We further analyze the overhead of OSPD along with its breakdown, to demonstrate our design is computationally efficient.
Overhead of Secure Partitioned Decoding.  Secure Partitioned Decoding introduces overhead due to (1) confidential computation (CC), (2) the absence of batch processing in CVMs, and (3) communication between CVMs and the LLM. For (1), Figure 11 illustrates the impact of confidential computing (CC) on the No protection baseline and SPD. It shows that CC results in approximately $15\%$ overhead compared to tranditional computing without CC. For (2) and (3), we present SPD’s latency breakdown in Figure 11. The six latency sources align with the processing steps in Figure 3. The QKV projection, public attention computation and MLP are common in traditional decoding. However, the unique overhead introduced by SPD stems from the other three sources. Notably, the “Receiving private attention” latency includes both the time the cloud LLM spends waiting for private attention computation to finish and the communication cost. This encapsulates the overhead from (2) and (3), as either the absence of batch processing or high communication costs can cause the LLM to remain idle. As shown in Figure 11, the latency of (2) and (3) (“Sending $Q_{new}$” and “Receiving private attention”) dominates the overhead, accounting for up to $40\%$ of the total latency. In contrast, the overhead from merging attentions is relatively minor.
Overhead of Prompt Obfuscation.  Prompt Obfuscation introduces overhead due to (1) sampling fake n-grams for virtual prompts, which increases TTFT, and (2) decoding for $\lambda$ virtual prompts. For (1), Figure 13 illustrates the sampling latency for virtual prompts. The average sampling time for an eight-token replacement using $\lambda=512$ and $\epsilon=1/32$ is around 1 second. Sampling multiple subsequences can be batched, for example, 16 subsequences take about 1.5 seconds. Lower $\epsilon$ results in faster sampling due to a reduced $\lambda$. For (2), the impact of decoding virtual prompts on latency is shown in Figure 13. When $\lambda\leq 8$, the overhead is less than $3\%$. However, the overhead increases as $\lambda$ grows: at $\lambda=64$ and $\lambda=128$, the overhead is approximately $7\%$ and $30\%$, respectively. This overhead also scales with the number of users. As shown in Figure 9, for 8 users, the overhead of $\lambda=8$ and $\lambda=16$ is around $8\%$ and $25\%$, respectively. Fundamentally, the overhead grows with the total number of prompts increases, as the LLM processes user prompts and virtual prompts equally. On the other hand, virtual prompts impose minimal overhead on the CVM memory footprint, thanks to our memory-efficient approach for storing virtual prompts (§5.1). In summary, the use of PO and its associated overhead reflects a key trade-off in our design, as outlined in §3.1, that is balancing prompt confidentiality (§5.2) with compute efficiency.
@@ 8 Concluding Remarks
Cloud-based LLM services are becoming pervasive in our daily lives, but raise privacy concerns since users must submit their prompts to the cloud. OSPD combines two techniques to protect user prompts from LLM providers: secure partitioned decoding (SPD) and prompt obfuscation (PO). SPD fully leverages the confidential computing capabilities of modern hardware for efficient and scalable confidential prompting. PO, on the other hand, leverages the statistical properties of the LLM itself to achieve confidentiality without encryption. Our proposed techniques have the potential to enable privacy-preserving LLM applications such as chatbots and AI assistants that interact with a large number of end users, without high operational costs.
Appendix A Appendix
A.1 Proof of Theorem 4.1
Let $Q\in\mathbb{R}^{d}$ be the query vector. Partition the key and value matrices $K\in\mathbb{R}^{n\times d}$ and $V\in\mathbb{R}^{n\times d}$ into private and public components:
$K=\begin{bmatrix}K_{\mathrm{Pvt}}\\ K_{\mathrm{Pub}}\end{bmatrix},\quad V=\begin{bmatrix}V_{\mathrm{Pvt}}\\ V_{\mathrm{Pub}}\end{bmatrix}.$
Compute the attention scores $s$ by:
$s=QK^{\top}=\begin{bmatrix}QK_{\mathrm{Pvt}}^{\top}&QK_{\mathrm{Pub}}^{\top}\end{bmatrix}=\begin{bmatrix}s_{\mathrm{Pvt}}&s_{\mathrm{Pub}}\end{bmatrix},$
where $s_{\mathrm{Pvt}}=QK_{\mathrm{Pvt}}^{\top}$ and $s_{\mathrm{Pub}}=QK_{\mathrm{Pub}}^{\top}$. Define the softmax denominators:
$\gamma=\sum_{i=1}^{n}\exp(s_{i})=\gamma_{\mathrm{Pvt}}+\gamma_{\mathrm{Pub}},$
with
$\gamma_{\mathrm{Pvt}}=\sum_{i=1}^{n_{\mathrm{Pvt}}}\exp(s_{\mathrm{Pvt},i}),\quad\gamma_{\mathrm{Pub}}=\sum_{i=1}^{n_{\mathrm{Pub}}}\exp(s_{\mathrm{Pub},i}).$
The attention output is:
$\sigma(s)^{\top}V$ $=\sum_{i=1}^{n}\frac{\exp(s_{i})}{\gamma}V_{i}$
$=\frac{1}{\gamma}\left(\sum_{i=1}^{n_{\mathrm{Pvt}}}\exp(s_{\mathrm{Pvt},i})V_{\mathrm{Pvt},i}+\sum_{i=1}^{n_{\mathrm{Pub}}}\exp(s_{\mathrm{Pub},i})V_{\mathrm{Pub},i}\right)$
$=\frac{\gamma_{\mathrm{Pvt}}}{\gamma}\left(\frac{1}{\gamma_{\mathrm{Pvt}}}\sum_{i=1}^{n_{\mathrm{Pvt}}}\exp(s_{\mathrm{Pvt},i})V_{\mathrm{Pvt},i}\right)$
$\quad+\frac{\gamma_{\mathrm{Pub}}}{\gamma}\left(\frac{1}{\gamma_{\mathrm{Pub}}}\sum_{i=1}^{n_{\mathrm{Pub}}}\exp(s_{\mathrm{Pub},i})V_{\mathrm{Pub},i}\right)$
$=\frac{\gamma_{\mathrm{Pvt}}}{\gamma}\left(\sigma(s_{\mathrm{Pvt}})^{\top}V_{\mathrm{Pvt}}\right)+\frac{\gamma_{\mathrm{Pub}}}{\gamma}\left(\sigma(s_{\mathrm{Pub}})^{\top}V_{\mathrm{Pub}}\right).$ (8)
Thus,
$\quad+\frac{\gamma_{\mathrm{Pub}}}{\gamma_{\mathrm{Pvt}}+\gamma_{\mathrm{Pub}}}\,\sigma(QK_{\mathrm{Pub}}^{\top})V_{\mathrm{Pub}},$ (9)
which completes the proof.
A.2 Proof of Theorem 5.2
To prove Theorem 5.2, we show that the two sets of subsequences are $\epsilon$-close over $\mathcal{D}_{\text{LM}}$ with Lemma A.1, and $(\epsilon+2\Delta)$-close over $\mathcal{D}$.
Lemma A.1.
Let $a_{1},a_{2},\ldots,a_{n}$ and $b_{1},b_{2},\ldots,b_{n}$ be sequences where $0\leq a_{i},b_{i}\leq 1$. If $|a_{i}-b_{i}|<\epsilon/n$ for all $i$, then $|\prod_{i=1}^{n}a_{i}-\prod_{i=1}^{n}b_{i}|<\epsilon$.
Part I  Let us denote each token in the subsequence $S_{\text{target}}$ as random variables $X_{1},\ldots,X_{n}$, where $X_{1}$ is the start token and $X_{n}$ is the end token. Let $f(X_{i})=P(X_{i}\mid X_{1},\ldots,X_{i-1})$ be the conditional probability of token $X_{i}$, which is modeled by the language model $LM$, given the previous tokens and the context $S$. Similarly, we represent the tokens in a fake sequence sampled by GQS as $Y_{1},\ldots,Y_{n}$.
GQS partitions the next token distribution into $\frac{1}{\epsilon^{\prime}}$ bins, where $\epsilon^{\prime}=\epsilon/n$, and each bin contains tokens with probabilities within $[m\epsilon^{\prime},(m+1)\epsilon^{\prime}]$, where $m$ is the bin index. Since GQS samples $Y_{i}$ from the same bin as $X_{i}$, it is true that $|f(X_{i})-f(Y_{i})|<\epsilon^{\prime}$ for all $i$.
Finally, we can bound the difference between the probability of the target sequence, $\prod_{i=1}^{n}f(X_{i})$, and the probability of the fake sequence, $\prod_{i=1}^{n}f(Y_{i})$, as $n\epsilon^{\prime}=\epsilon$, using Lemma A.1. That is,
$|\prod_{i=1}^{n}f(X_{i})-\prod_{i=1}^{n}f(Y_{i})|<\epsilon$
This is true for every pair of target and fake subsequences. Based on Definition 5.1, since the maximum difference between the probabilities (modeled by $\mathcal{D}_{\text{LM}}$) of any target and fake subsequences is $\epsilon$, the two sets of subsequence are $\epsilon$-close over $\mathcal{D}_{\text{LM}}$.
Part II  Let us denote the subsequence $S_{target}$ as $s$, and a fake n-gram sampled by GQS as $s^{\prime}$. According to Part I and the definitions in §5.2, we have
$|P_{\mathcal{D}}(s)-P_{\mathcal{D}}(s^{\prime})|$ $=|P_{\mathcal{D}}(s)-P_{\mathcal{D}_{\text{LM}}}(s)+P_{\mathcal{D}_{\text{LM}}}(s)-P_{\mathcal{D}_{\text{LM}}}(s^{\prime})$
$\quad+P_{\mathcal{D}_{\text{LM}}}(s^{\prime})-P_{\mathcal{D}}(s^{\prime})|$
$\leq|P_{\mathcal{D}}(s)-P_{\mathcal{D}_{\text{LM}}}(s)|+|P_{\mathcal{D}_{\text{LM}}}(s)-P_{\mathcal{D}_{\text{LM}}}(s^{\prime})|$
$\quad+|P_{\mathcal{D}_{\text{LM}}}(s^{\prime})-P_{\mathcal{D}}(s^{\prime})|$
$\leq\epsilon+2\Delta.$
This is true for every pair of target and fake subsequences. Based on Definition 5.1, since the maximum difference between the probabilities (modeled by $\mathcal{D}$) of any target and fake subsequences is $\epsilon+2\Delta$, the two sets of subsequences are $(\epsilon+2\Delta)$-close over $\mathcal{D}$.
A.3 Example and Dataset Preparation for PO
[AGE] 17-year-old [GENDER] male, has come to the [ORGANIZATION NAME] student health clinic complaining of [DISEASE] heart pounding. [PERSON NAME] Mrs. Loraine Wicks mother has given verbal consent for a history, physical examination, and treatment - began [DATES] 2-3 months ago, sudden, intermittent for [DATES] 2 days (lasting 3-4 min), worsening, non-allev/aggrav - associated with dyspnea on exertion and rest, stressed out about school - reports he feels like his heart is jumping out of his chest - ROS: denies chest pain, diaphoresis, weight loss, chills, fever, nausea, vomiting, pedal edema - PMH: none, meds: Adderall (from a friend), NKDA - FH: father had MI recently, mother has thyroid disease - SH: non-smoker, marijuana [DATES] 5-6 months ago, 3 beers on the weekend, basketball at school - SH: no STD.
Figure 15: Example entry from the clinical dataset. Yellow-masked text indicates redacted information, with brackets denoting its PII category.
Resume Wizard [PERSON NAME] Jane Doe E-Mail: [EMAIL] jane.doe@example.com Phone: [ID] (123) 456-7890 (M) [ORGANIZATION NAME] Finance & Accounts, Costing Profile Introduction A dynamic professional with a qualitative experience of [AGE] 3 years 2 months in the areas of Finance & Accounts, Product Costing & MIS Reporting. Presently working as PROCESS LEAD at [ORGANIZATION NAME] Tech Solutions Inc. Previously worked as LEAD F&A OPERATIONS at [ORGANIZATION NAME] Global Tech with [ORGANIZATION NAME] Soft Drinks Co Project from [DATES] 1st April 2018 to [DATES] 30th June 2019. Involved in the standardization of process & MIS reporting files and contributed a higher rate of organic growth. An effective communicator with excellent relationship-building & interpersonal skills. Strong analytical, problem-solving & organizational abilities. Areas of interest include budgeting & preparation of AOP (Annual Operating Plans). Familiar with SAP FICO modules, MS Office tools.
Figure 16: Example entry from the resume dataset. Note that we anonymized this example with placeholders such as “Jane Doe.” and “(123) 456-7890” for privacy.
For the evaluation of PO, we selected the initial 500 entries from clinical dialogue datasets and 1500 entries from USMLE Step 2 passages. Since the initial datasets were not tagged with PII information, we used ChatGPT to automatically tag HIPAA-sensitive information according to the guidelines from [20], along with its PII category. Each entry in these datasets averaged 13 redacted PII segments. An example of a dataset item is shown in Figure 15. For the resume dataset, we extracted the first 1500 entries and used Google Cloud’s PII filtering tool to annotate personal information, resulting in an average of 18 PII segments per entry. An anonymized example is presented in Figure 16.
Example Output from GQS.  For demonstration, we present the result of running GQS on the name “Mrs. Loraine Wicks” in Figure 15, sampled with $\epsilon=1/32$ and $\lambda_{\text{max}}=512$. Due to space constraints, only the first 30 outputs are shown.
  1. 1.
    Mrs. Loraine Wicks
  2. 2.
    Mrs. Lillian Brown
  3. 3.
    Mrs. Loraine Park
  4. 4.
    Mrs. Lydia Jackson
  5. 5.
    Mrs. Lydia Williams
  6. 6.
    Mrs. Lillian Thomas
  7. 7.
    Mrs. Wilson
  8. 8.
    Mrs. Lydia Grant
  9. 9.
    Mrs. Lillian Taylor
  10. 10.
    Mrs. Lydia Green
  11. 11.
    Mrs. Loraine Warner
  12. 12.
    Mrs. Lydia Martin
  13. 13.
    Mrs. Loraine Thompson
  14. 14.
    Mrs. Loraine Wood
  15. 15.
    Mrs. Loraine Brown
  16. 16.
    Mrs. Lillian Grey
  17. 17.
    Mrs. Lillian Jenkins
  18. 18.
    Mrs. Lillian Watson
  19. 19.
    Mrs. Lillian Anderson
  20. 20.
    Mrs. Loraine Williams
  21. 21.
    Mrs. Lillian Winter
  22. 22.
    Mrs. Lillian Abraham
  23. 23.
    Mrs. Lillian Cooper
  24. 24.
    Mrs. Lydia Davis
  25. 25.
    Mrs. Lillian Kim
  26. 26.
    Mrs. Lillian Thompson
  27. 27.
    Mrs. Lydia White
  28. 28.
    Mrs. Lillian Davis
  29. 29.
    Mrs. Lillian Carter
  30. 30.
    Mrs. Loraine Miller
A.4 Limitation of OSPD
Here, we discuss some limitations in OSPD as well as new opportunities.